OpenStudy (anonymous):

Write the given expression as a single logarithm. -3ln(x)+ln(4y)-ln(3z)

OpenStudy (mertsj):

ln(x^-3)(4y)(3z)= ln(12yz/x^3)

OpenStudy (anonymous):

thats wrong for some reason

OpenStudy (anonymous):

I believe you can write ln(x^-3*4y/3z)

OpenStudy (anonymous):

\[-3\ln x+\ln4y-\ln3z\]\[\ln x^{-3}+\ln4y-\ln3z\]\[\ln 4yx^{-3}-\ln3z\]\[\ln \frac{4yx^{-3}}{3z}\]

OpenStudy (anonymous):

Thanks guys! Since you guys are familiar with all this, can you help me with the other problems I posted?